OpenStudy (anonymous):

f(x) = x^3 + ax^2 + bx + c Find the consonants a,b, and c so that the graph will increase to the point (-3,18), decrease to the point (1, -14), and then continue increasing.

myininaya (myininaya):

So we have critical numbers do you agree? What are the 2 critical numbers?

myininaya (myininaya):

notisaiah?

OpenStudy (phi):

The critical points are found by taking the first derivative.

myininaya (myininaya):

Great!

myininaya (myininaya):

Now since these are critical numbers and this is a poly that would mean f'(-3)=0 and f'(1)=0

OpenStudy (anonymous):

okay

myininaya (myininaya):

So now phi said you need to find the first derivative and he is correct

myininaya (myininaya):

Could you tell me what you get for f'(x)?

OpenStudy (anonymous):

Wait, so I find the derivative of the function, then i plug in -3 and 1 ?

OpenStudy (anonymous):

f`(x) = 3x^2 + 2ax + b

myininaya (myininaya):

Great! So Evaluate f' at x=-3 and evaluate f' at x=1 Do we not have more information?

myininaya (myininaya):

I'm not seeing how to find c right now but this will give you two linear equations that you can solve for a and b

OpenStudy (anonymous):

No, that is it I get f`(-3) = 27 - 6a + b and f`(1) = 3 + 2a + b

myininaya (myininaya):

And both of those are =0

myininaya (myininaya):

27-6a+b=0 3+2a+b=0

OpenStudy (anonymous):

yes.

myininaya (myininaya):

This was the set of linear equations I was talking about

OpenStudy (anonymous):

right.

myininaya (myininaya):

Oh I know how to get c I'm an idiot I wasn't thinking that far ahead but we are given two points on f

myininaya (myininaya):

So once you find a and b from there We will go back to the original function f and do f(-3)=18 and f(1)=-14 to find c

myininaya (myininaya):

or*

OpenStudy (anonymous):

could you elaborate please

myininaya (myininaya):

Which part? Finding a and b? Or using f(1)=-14 to find c?

OpenStudy (anonymous):

What I have right now are the two functions I get f`(-3) = 27 - 6a + b + c and f`(1) = 3 + 2a + b + c

OpenStudy (anonymous):

oops, without the c okay let me find the values for a and b

myininaya (myininaya):

Remember those are both equal to 0

OpenStudy (phi):

remember these points are at a max or a min, so the slope is 0

OpenStudy (anonymous):

okay so I have the two functions f`(-3) = 27 - 6a + b and f`(1) = 3 + 2a + b where b = -9 , a = 3

OpenStudy (phi):

now use the original equation with your a,b values

OpenStudy (anonymous):

So now I have f`(x) = x^3 + 3x^2 - 9x + c How do I find the c value

OpenStudy (phi):

It's f(x) not f'(x). f(x) is another name for the 2nd value (y value) in an (x,y) pair that satisfies the equation. You have 2 points. Choose one of them.

OpenStudy (anonymous):

I see. thank you !